A cat weighs 50N. I lift it so that it accelerates at a rate of 2m/s2. How much lifting force am I applying to the cat?
1
Expert's answer
2020-04-13T10:20:52-0400

The mass of a car

"m=\frac{W}{g}=\frac{50\:\rm N}{10\:\rm N/kg}=5\:\rm kg."

The Newton's second law says

"ma=F_{\rm lift}-W."

Hence, the lift force

"F_{\rm lift}=ma+W=\\
=(5\:\rm kg)\times (2\:\rm m/s^2)+(50\:\rm N)=60\:\rm N."